A sample of 680 adults in the United States were asked the following question from the General Social Survey:  “Do you sometimes drink more than you think you should?” The table below shows their responses.  

Sex

Yes

No

Total

Men

151

177

328

Women

92

260

352

Total

243

437

680

  1. What is the percent probability that a randomly selected woman thinks she drank more than she should?
    1. 13.5%
    2. 26.1%
    3. 35.7%
    4. 37.9%
Answer #1

Solution:

Probability that a randomly selected woman thinks she drank more than she should

=P(Yes|Woman)

=\frac{P(Woman \cap Yes)}{P(Woman)}

=\frac{92/680}{352/680}

=\frac{92}{352}

=0.261

=26.1\%

Hence the correct option is (B) 26.1%.

**For further queries, please comment below.

Thank You.